What is the minimum distance required for the car to stop?Suppose that a 1.00x103 kg car is traveling at 12.5 m/s (~45 km/hr).  Its brakes can apply a force of 5.00x103 N.


1
Expert's answer
2021-05-18T17:29:52-0400
"v^2=2as\\v^2=2\frac{F}{m}s\\12.5^2=2\frac{5000}{1000}s\\s=15.6\ m"